# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 5.2. Existence of the mixed discrete conformal structure $\mathrm{I}$
There exists $j\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), such that $f_{j}^{(m)}\rightarrow+\infty$.
For case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law,
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
The case $\mathrm{(ii)}$ follows from Lemma 5.5 and the case $\mathrm{(iii)}$ follows from Lemma 5.6, which both contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. Q.E.D.
@@
Note that in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), there is no $j\in B$ such that $f_{j}^{(m)}\rightarrow-\infty$. Otherwise, $u_{j}=-e^{-f_{j}}\rightarrow-\infty$. According to the admissible space $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ defined by (44), the equation $u_{j}+u_{k}>-2\sqrt{\eta_{jk}}$ implies $u_{k}\rightarrow+\infty$, which contradicts with the fact that $u_{k}\in\mathbb{R}_{<0}$.
6. Rigidity and existence of the mixed discrete conformal structure II
In this section, we consider the rigidity and the image of $K$ for the mixed discrete conformal structure $\mathrm{II}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (5) and (9). As mentioned in Remark 1.8, if $d_{rs}<0$, then $l_{rs},l_{rt}$ are given by (9) and $l_{st}$ is given by (5), where $\{r,s,t\}=\{i,j,k\}$. Note that $\eta\in\mathbb{R}_{>0}^{E}$ by (5) and $\zeta\in\mathbb{R}^{E}$ by (9). To simplify the notation, we use a new parameter $\xi$ to replace $\eta$ and $\zeta$.
@@ 6. Rigidity and existence of the mixed discrete conformal structure II
Without loss of generality, we can always assume $d_{ij}<0$ and $d_{ik}<0$. Then for any right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij}$ and $l_{ik}$ are given by (9) and the edge length $l_{jk}$ is given by (5), i.e.,
(46) $\cosh l_{ij}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{jk}e^{f_{j}+f_{k}}.$
6.1. Admissible space of mixed discrete conformal structure $\mathrm{II}$
@@
The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46) is defined to be
$\Omega_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
@@
Since $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}=\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{k}e^{2f_{k}}}}$, $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{k}e^{2f_{k}}}}$, then by by the proof of Lemma 2.22, we have
@@
To obtain the admissible space $\mathcal{U}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ with respect to the discrete conformal factors $u$, we need to characterize $l_{ij}>0$ and $l_{jk}>0$. Since the case that $l_{jk}>0$ has been already discussed in Subsection 3.1, we just need to discuss the case $l_{ij}>0$. By Remark 2.23, we have
@@
If $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\cosh l_{ij}=1+\xi_{ij}e^{u_{j}-u_{i}}>1$, which implies $\xi_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
@@
If $\alpha_{i}=1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}}=\coth u_{i}+\xi_{ij}e^{u_{j}}\frac{1}{\sinh u_{i}}>1$, which implies $\xi_{ij}e^{u_{i}+u_{j}}+1>0$ by $u_{i}>0$. Thus if $\xi_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<0$, then $u_{i}+u_{j}<\log(-\frac{1}{\xi_{ij}})$. For simplicity, we only consider the case $\xi_{ij}\geq 0$ in the following.
@@
If $\alpha_{i}=-1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}}=\tanh u_{i}+\xi_{ij}e^{u_{j}}\frac{1}{\cosh u_{i}}>1$, which implies $\xi_{ij}e^{u_{i}+u_{j}}>1$. Thus $\xi_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$.
@@
If $\alpha_{i}=0$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{j}+\xi_{ij}e^{-u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $e^{u_{i}+u_{j}}+\xi_{ij}>0$ by $u_{j}<0$. Thus if $\xi_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<0$, then $u_{i}+u_{j}>\log(-\xi_{ij})$.
@@
If $\alpha_{i}=1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1+e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\coth u_{j}+\xi_{ij}\frac{1}{\sinh u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $\cosh(u_{i}+u_{j})+\xi_{ij}>0$ by $u_{i}>0$ and $u_{j}<0$. Thus if $\xi_{ij}>-1$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<-1$, then $u_{i}+u_{j}>\mathrm{arccosh}\ (-\xi_{ij})$ or $u_{i}+u_{j}<-\mathrm{arccosh}\ (-\xi_{ij})$. For simplicity, we only consider the case $\xi_{ij}>-1$ in the following.
@@
If $\alpha_{i}=-1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1+e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\coth u_{j}+\xi_{ij}\frac{1}{\cosh u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $\sinh(u_{i}+u_{j})+\xi_{ij}>0$ by $u_{j}<0$. Thus $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$.
@@
If $\alpha_{i}=0$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{j}+\xi_{ij}e^{-u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $e^{u_{i}+u_{j}}<\xi_{ij}$. Thus $\xi_{ij}>0$ and $u_{i}+u_{j}<\log\xi_{ij}$.
@@
If $\alpha_{i}=1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1-e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\tanh u_{j}+\xi_{ij}\frac{1}{\sinh u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $\sinh(u_{i}+u_{j})<\xi_{ij}$. Thus $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$.
@@
If $\alpha_{i}=-1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1-e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\tanh u_{j}+\xi_{ij}\frac{1}{\cosh u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $\cosh(u_{i}+u_{j})<\xi_{ij}$. Thus $\xi_{ij}>1$ and $-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij}$.
For simplicity, we use triples $(\alpha_{i},\alpha_{j},\alpha_{k})$ to represent different types of the admissible spaces of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46). By permutations and combinations, the admissible space $\Omega_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ of $f$ can be transferred to 27 types of the admissible spaces of $u$. Note that the positions of $j$ and $k$ are inter-changeable in (46), we regard $(\alpha_{i},\alpha_{j},\alpha_{k})$ and $(\alpha_{i},\alpha_{k},\alpha_{j})$ as the same type. Therefore, there exist 18 types of the admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ of $u$, as shown below. Here $n=1,2,...,18$. For simplicity, we write $\mathcal{U}^{(n)}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ as $\mathcal{U}^{(n)}_{ijk}(\xi)$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,0)$. Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1. By the above discussion, we have $\xi_{ij}>0,\ (u_{j}-u_{i})\in\mathbb{R}$ and $\xi_{ik}>0,\ (u_{k}-u_{i})\in\mathbb{R}$. Therefore, the admissible space of $u$ is defined as
(48) $\mathcal{U}^{(1)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{3}|u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.

### v3

@@ 6.2. Existence of the mixed discrete conformal structure III
There exists $j\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48) or (7), such that $f_{j}^{(m)}\rightarrow+\infty$.
The proof for the case (i) is identical to that of Theorem 3.5 and is thus omitted. For the case (ii), by Lemma 6.8, each generalized angle incident to the boundary component $i\in B$ converges to 0. Consequently, $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=0$, which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. For the case (iii), if all right-angled hyperbolic hexagons are Type-III, then Lemma 6.9 implies $\lim_{m\rightarrow+\infty}K_{j}^{(m)}=0$; if not all right-angled hyperbolic hexagons are Type-III, then by Lemma 6.7 and Lemma 6.9, we have $\lim_{m\rightarrow+\infty}K_{j}^{(m)}=0$. Both scenarios contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. Q.E.D.
@@
Note that in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), there is no $j\in B$ such that $f_{j}^{(m)}\rightarrow-\infty$. Suppose, to the contrary, that such a $j$ exists, then $u_{j}=-e^{-f_{j}}\rightarrow-\infty$. By the definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ in (49), the inequality $u_{j}+u_{k}>-2\sqrt{\eta_{jk}}$ implies $u_{k}\rightarrow+\infty$, which contradicts the fact that $u_{k}\in\mathbb{R}_{<0}$. An analogous argument applies to a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (7).
7. Rigidity of the mixed discrete conformal structure II
In this section, we investigate the rigidity of the mixed discrete conformal structure II, specifically, an ideally triangulated surface with boundary admitting both discrete conformal structures (A2) and (B2). All right-angled hyperbolic hexagons are either Type-II or a combination of Type-II right-angled hyperbolic hexagons and those with edge lengths given by (6). Here, a right-angled hyperbolic hexagon is Type-II if its edge lengths are given by both (6) and (9). By Remark 1.6, we have $\alpha\equiv-1$ for both (6) and (9).
@@ 7. Rigidity of the mixed discrete conformal structure II
As noted in Remark 1.7, if $d_{rs}<0$, then $d_{rt}<0,\ d_{sr}>0,\ d_{tr}>0,\ d_{st}>0,\ d_{ts}>0$, and $l_{rs},l_{rt}$ are given by (9) and $l_{st}$ is given by (6), where $\{r,s,t\}=\{i,j,k\}$. Without loss of generality, we assume $d_{ij}<0$ in this section. This implies $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$. For the Type-II right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (9), and $l_{jk}$ is given by (6), namely,
(52) $\cosh l_{ij}$ $=-\sqrt{(e^{2f_{i}}-1)(e^{2f_{j}}-1)}+\eta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=-\sqrt{(e^{2f_{i}}-1)(e^{2f_{k}}-1)}+\eta_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=\sqrt{(e^{2f_{j}}-1)(e^{2f_{k}}-1)}+\eta_{jk}e^{f_{j}+f_{k}}.$
Here $\eta_{ij}\in\mathbb{R}_{>0}$, $\eta_{ik}\in\mathbb{R}_{>0}$, and $\eta_{jk}\in(-1,+\infty)$.
@@
For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (52), the admissible space of the discrete conformal factor $f$ is defined as
$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}_{>0}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
By combining (30) and Remark 2.22 (iv), we consider $u\in(-\frac{\pi}{2},0)^{N}$ for simplicity. It follows that $\cosh l_{ij}=-(-\tan u_{i})\frac{-1}{\tan u_{j}}+\eta_{ij}\frac{1}{\cos u_{i}}\frac{-1}{\sin u_{j}}>1$, which implies $\sin(u_{i}+u_{j})>-\eta_{ij}$. Consequently, $u_{i}+u_{j}\in(-\pi,0)$ for $\eta_{ij}\in[1,+\infty)$, and $u_{i}+u_{j}>-\arcsin\eta_{ij}$ for $\eta_{ij}\in[0,1)$. Moreover, from (45), $\cosh l_{jk}>1$ implies $u_{j}+u_{k}\in(-\pi,0)$ for $\eta_{jk}\in[1,+\infty)$ and $u_{j}+u_{k}>-\arccos\eta_{jk}$ for $\eta_{jk}\in(-1,1)$.
Theorem 7.1.
Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (52), and the weight $\eta_{ij}\in[0,+\infty)$, $\eta_{ik}\in[0,+\infty)$, $\eta_{jk}\in(-1,+\infty)$. Then the admissible space
[…]
$\sin(u_{i}+u_{k})>-\eta_{ik},\cos(u_{j}+u_{k})>-\eta_{jk}\}.$
of $u$ is a convex polytope.
Corollary 7.2.
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary. If one of the following conditions holds:
(1):
All right-angled hyperbolic hexagons are Type-II (i.e., the edge lengths of any right-angled hyperbolic hexagons are given by (52)), and the weight $\eta$ satisfies the condition in Theorem 7.1;
(2):
Not all right-angled hyperbolic hexagons are Type-II (i.e., there exist some right-angled hyperbolic hexagons with edge lengths given by (6)). In this case, the weight $\eta$ satisfies the condition in Theorem 7.1 for those right-angled hyperbolic hexagons with edge lengths given by (52), and $\eta\in(-1,+\infty)^{E}$ for those with edge lengths given by (6).
Then the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
Proof.
The case (1) follows from Theorem 7.1, and the case (2) follows from (46) and Theorem 7.1. Q.E.D.
Theorem 7.3.
Under the same assumptions as those in Theorem 7.1, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite on $\mathcal{U}_{ijk}(\eta)$.
Proof.
For simplicity, we adopt the notations introduced in Subsection 3.2. Define
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-\sqrt{e^{2f_{i}}-1}&0&0\\ 0&\sqrt{e^{2f_{j}}-1}&0\\ 0&0&\sqrt{e^{2f_{k}}-1}\\ \end{array}\right):=Q_{7}.$
Then $\Lambda_{ijk}$ can be expressed as
$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{7},$
where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively. Furthermore, $\det Q_{1}>0$. For any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (52), since $d_{ij}<0$, $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$, it follows from (39) that $\det Q_{2}<0$. The remaining part of the proof is analogously to that of Theorem 6.3, and is thus omitted here.
Q.E.D.
Combining Theorem 4.1 and Theorem 7.3, we obtain the following result.
Corollary 7.4.
Under the same assumptions as those in Corollary 7.2, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
The following theorem establishes the rigidity of the mixed discrete conformal structure II, which generalizes Theorem 1.8 (v). As its proof is analogous to that of Theorem 3.4, it is omitted here.
Theorem 7.5.
Under the same assumptions as those in Corollary 7.2, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
Remark 7.6.
The existence of the mixed discrete conformal structure II cannot be established. Specifically, the boundary of the admissible space $\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ in $[-\infty,+\infty]^{N}$ include the case that there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow 0^{+}$. For this case, we are unable to prove the closedness of the curvature map.
8. Rigidity and existence of the mixed discrete conformal structure I
In this section, we investigate the rigidity and existence of the mixed discrete conformal structure I, specifically, an ideally triangulated surface with boundary admitting both discrete conformal structures (A1) and (B1). All right-angled hyperbolic hexagons are either Type-I or a combination of Type-I right-angled hyperbolic hexagons and those with edge lengths given by (5). Here, a right-angled hyperbolic hexagon is Type-I if its edge lengths are given by both (5) and (8).
As noted in Remark 1.7, if $d_{rs}<0$, then $d_{rt}<0,\ d_{sr}>0,\ d_{tr}>0,\ d_{st}>0,\ d_{ts}>0$, and $l_{rs},l_{rt}$ are given by (8) and $l_{st}$ is given by (5), where $\{r,s,t\}=\{i,j,k\}$. Without loss of generality, we assume $d_{ij}<0$ in this section. This implies $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$. For the Type-I right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (8), and $l_{jk}$ is given by (5), namely,
(53) $\cosh l_{ij}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{k}e^{2f_{k}})}+\eta_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\eta_{jk}e^{f_{j}+f_{k}}.$
8.1. Admissible space of the mixed discrete conformal structure I
For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (53), the admissible space of the discrete conformal factor $f$ is defined as
$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
@@ 8.1. Admissible space of the mixed discrete conformal structure I
Since $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}=\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{k}e^{2f_{k}}}}$, and $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{k}e^{2f_{k}}}}$, it follows from the proof of Lemma 2.21 that
@@
To characterizes the admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$, it is necessary to derive the conditions under which the edge lengths $l_{ij}>0$ and $l_{jk}>0$. Since the case that $l_{jk}>0$ has already been addressed in Subsection 3.1, we focus on analyzing $l_{ij}>0$ here. By Remark 2.22, the following results are obtained:
@@
If $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\cosh l_{ij}=1+\eta_{ij}e^{u_{j}-u_{i}}>1$. This implies $\eta_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
@@
If $\alpha_{i}=1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}}=\coth u_{i}+\eta_{ij}e^{u_{j}}\frac{1}{\sinh u_{i}}>1$. Since $u_{i}>0$, it follows that $\eta_{ij}e^{u_{i}+u_{j}}+1>0$. Consequently, if $\eta_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\eta_{ij}<0$, then $u_{i}+u_{j}<\log(-\frac{1}{\eta_{ij}})$. For simplicity, we only consider the case $\eta_{ij}\geq 0$ in the following.
@@
If $\alpha_{i}=-1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}}=\tanh u_{i}+\eta_{ij}e^{u_{j}}\frac{1}{\cosh u_{i}}>1$. This implies $\eta_{ij}e^{u_{i}+u_{j}}>1$, so $\eta_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$.
@@
If $\alpha_{i}=0$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{j}}}+\eta_{ij}e^{f_{i}+f_{j}}=-\coth u_{j}+\eta_{ij}e^{-u_{i}}\frac{-1}{\sinh u_{j}}>1$. Since $u_{j}<0$, it follows that $e^{u_{i}+u_{j}}+\eta_{ij}>0$. Consequently, if $\eta_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\eta_{ij}<0$, then $u_{i}+u_{j}>\log(-\eta_{ij})$.
@@
If $\alpha_{i}=1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1+e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\coth u_{j}+\eta_{ij}\frac{1}{\sinh u_{i}}\frac{-1}{\sinh u_{j}}>1$. Since $u_{i}>0$ and $u_{j}<0$, it follows that $\cosh(u_{i}+u_{j})+\eta_{ij}>0$. Consequently, if $\eta_{ij}>-1$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\eta_{ij}<-1$, then $u_{i}+u_{j}>\mathrm{arccosh}\,(-\eta_{ij})$ or $u_{i}+u_{j}<-\mathrm{arccosh}\,(-\eta_{ij})$. For simplicity, we only consider the case $\eta_{ij}>-1$ in the following.
@@
If $\alpha_{i}=-1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1+e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\coth u_{j}+\eta_{ij}\frac{1}{\cosh u_{i}}\frac{-1}{\sinh u_{j}}>1$. Since $u_{j}<0$, it follows that $\sinh(u_{i}+u_{j})+\eta_{ij}>0$. Thus $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\eta_{ij})$.
@@
If $\alpha_{i}=0$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{j}}}+\eta_{ij}e^{f_{i}+f_{j}}=-\tanh u_{j}+\eta_{ij}e^{-u_{i}}\frac{1}{\cosh u_{j}}>1$. This implies $e^{u_{i}+u_{j}}<\eta_{ij}$, so $\eta_{ij}>0$ and $u_{i}+u_{j}<\log\eta_{ij}$.
@@
If $\alpha_{i}=1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1-e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\tanh u_{j}+\eta_{ij}\frac{1}{\sinh u_{i}}\frac{1}{\cosh u_{j}}>1$. This implies $\sinh(u_{i}+u_{j})<\eta_{ij}$, so $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij}$.
@@
If $\alpha_{i}=-1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1-e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\tanh u_{j}+\eta_{ij}\frac{1}{\cosh u_{i}}\frac{1}{\cosh u_{j}}>1$. This implies $\cosh(u_{i}+u_{j})<\eta_{ij}$, so $\eta_{ij}>1$ and $-\mathrm{arccosh}\,\eta_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\,\eta_{ij}$.
For simplicity, we adopt the triple notation $(\alpha_{i},\alpha_{j},\alpha_{k})$ to classify the distinct types of admissible spaces of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (53). Through permutations and combinations, the admissible space $\Omega_{ijk}(\eta)$ of $f$ can be transformed into 27 distinct types of admissible spaces of $u$. However, due to the interchangeability of the positions of $j$ and $k$ in (53), the triples $(\alpha_{i},\alpha_{j},\alpha_{k})$ and $(\alpha_{i},\alpha_{k},\alpha_{j})$ are considered equivalent and thus belong to the same type. This reduction results in a total of 18 distinct types of admissible spaces of $u$, denoted by $\mathcal{U}^{(n)}_{ijk}(\eta)$, where $n=1,2,...,18$. These types are enumerated as follows.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,0)$. In Subsection 3.1, it follows that $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$ and $\eta_{jk}>0$. Building upon the preceding analysis, we deduce $\eta_{ij}>0$ with $(u_{j}-u_{i})\in\mathbb{R}$, and $\eta_{ik}>0$ with $(u_{k}-u_{i})\in\mathbb{R}$. Consequently, the admissible space of $u$ is defined as
(55) $\mathcal{U}^{(1)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{3}\mid u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$.